A piece of newly purchased industrial equipment costs $977,000 and is classified as seven-year property under MACRS. The MACRS depreciation schedule is shown in Table 10.7. Calculate the annual depreciation allowances and end-of-the-year book values for this equipment.
Answer #1

Compute the annual depreciation and end of the year book value as follows of the End of the year book value C-E Cost of Begin
Answer #2

Calculation of Depreciation using MACRS Schedule :

Year Depreciation Calculation Depreciation End of the year book value
Year 1 =977000*14.29% 139613.3 837386.7
Year 2 = 977000*24.49 239267.3 737732.7
Year 3 =977000*17.49 170877.3 806122.7
Year 4 =977000*12.49 122027.3 854972.7
Year 5 =977000*8.93 87246.1 889753.9
Year 6 =977000*8.92 87148.4 889851.6
Year 7 =977000*8.93 87246.1 889753.9